f 5 pupils are seated on each bench in a classroom, 4 will be left without a place. If 6 pupils are seated on each bench, 2 places remain on one bench. How many pupils and how many benches are there?

Answer by ewatrrr(24785)

 
Hi,
Let x represent the number of benches
5x + 4 = 6x-2
6 = x, the number of benches. 34 Students
CHECKING our answer***
30+4 = 36-2 = 34
Wish You the Best in your Studies.